OpenStudy (anonymous):

why f(x) = x^3 doesn't have relative extrema at 0

OpenStudy (anonymous):

f'(x)=3x^2 first derivative test...now let's test some values around the critical point f'(x)=0 3x^2=0 x=0 \[(-\infty, 0), (0, \infty)\] f'(1)=3 f'(-1)=3 No sign change here, no relative extrema

OpenStudy (anonymous):

whereas f(x) = x^2 have rel extrema at 0

OpenStudy (anonymous):

got it thanxxxxxxxx

OpenStudy (anonymous):

f(x)=x^2 f'(x)=2x 2x=0 critical point x=0 \[(-\infty, 0), (0, \infty)\] f'(-1)=-2 f'(1)=2 negative to positive ==>> local min

myininaya (myininaya):

Increasing to Decreasing menas we have a local max Decreasing to Increasing measn we have a local min In other words you have to have a switch! f=x^3 is increasing from (-inf,inf) there is no local extrema(or abs extrema) f=x^2 has local min since we have the switch is from decreasing to increasing

myininaya (myininaya):

on (-inf,inf)*